OpenStudy (anonymous):

Solve for x: 4|2x − 2| + 10 = 34.

pooja195 (pooja195):

Add -10 to both sides. 4(|2x−2|)+10+−10=34+−10 4(|2x−2|)=24 Step 2: Divide both sides by 4. (|2x−2|)/4=24/4 |2x−2|=6 Now set up 2 equations with a pos and neg solution 2x-2=-6 2x-2=6 Then for both equations solve for x you will have TWO answers

OpenStudy (anonymous):

thank u can u help me with another one @pooja195

pooja195 (pooja195):

sure :)

OpenStudy (anonymous):

An inequality is shown below: −np − 5 ≤ 4(c − 2) Which of the following solves for n?

pooja195 (pooja195):

what are the choices?

OpenStudy (anonymous):

nvm my parents helped me out but can u help me wit other ones if your not busy

pooja195 (pooja195):

How many more?

OpenStudy (anonymous):

depends on how many are hard

pooja195 (pooja195):

.-. ok repost